Out of 30 customers at a coffee shop. 14 buy a coffee only,6 buy a tea only and 8 order a doughnut only. State P(coffee order/order doughnut)
1
Expert's answer
2018-05-10T10:08:08-0400
Out of 30 customers at a coffee shop. 14 buy a coffee only,6 buy a tea only and 8 order a doughnut only.
Assume that events 'coffee order' and 'order doughnut' are independent.
Then
P(coffee order/order doughnut) = P(coffee order)=14/30=7/15.